OpenStudy (anonymous):

When the sum of a number plus 3 is squared, it is 9 more than the sum of the number plus 2 when squared. @alli14344

OpenStudy (anonymous):

(x+3)^2 = 9+(x+2)^2 x^2 + 6x + 9 = 9 + x^2 + 4x +4 6x = 4x + 4 -2x = 4 x = -2

OpenStudy (anonymous):

:)